Consider An Economy In Which The Consumption Function Takes The Following Simple Algebraic Form: C = 300+0.75DI And In Which Investment (I) Is Always $900 And Net Exports Are Always -$100. Government Purchases Are Fixed At $1,300 And Taxes Are Fixed At $1,200. Find The Equilibrium Level Of GDP.
1
Expert's answer
2020-11-09T07:00:50-0500

C = 300 + 0.75DI (Given)


Since DI = Y - T = Y -1,200, Substituting DI = Y - 1200


C = 300 + 0.75(Y – 1200)


C = 300 + 0.75Y – 900


C = -600 + 0.75Y


The Composition of equilibrium GDP


Y = C + I + G + (X – IM)


Since, I = $900, G = $1,300, (X - IM) = -100, Substituting these given values in the equation, we get


Y = -600 + 0.75Y + 900 + 1300 – 100


Y = 0.75Y + 1500


0.25Y= 1500


Y = (1/0.25) * 1500


Y = 4 * 1500 = 6000


Hence the equilibrium level of GDP is $ 6000